OpenStudy (anonymous):

Hank bought a new graduated cylinder for his chemistry class. It holds 973.36 mL of liquid. If the cylinder has a radius of 4.6 cm, then how tall is the cylinder? (1 milliliter is equal to 1 cubic centimeter) A. 44,775 cm B. 106 cm C. 46 cm D. 212 cm

OpenStudy (danjs):

cylinder Volume V = pi * radius^2 * height put in V and radius, and solve

OpenStudy (danjs):

ml and cm^3 are the same here

OpenStudy (anonymous):

Its asking for the height...not volume

OpenStudy (danjs):

yes, Volume = pi * radius^2 * height put in the given volume and radius, solve for the height

OpenStudy (anonymous):

So .....H=V/r2...right?

OpenStudy (danjs):

\[ h =\frac{ V }{ \pi*r^2 }\]

OpenStudy (anonymous):

Oh yeah thank you ^~^

OpenStudy (danjs):

=973.36 / (3.14*4.6*4.6) =973.36/66.44 =14.6 answer is not there, if you leave the pi out it is 46, maybe they made a mistake

OpenStudy (anonymous):

Nope I got it ^~^

OpenStudy (anonymous):

It left the pi out so it was 46 XP